a cargo ship with a mass of 5×105 kg floating in water carrying loads with unknown weight, WL. Suppose the cargo ship displaces an amount of 6300 m3 seawater while carrying the loads, determine the maximum weight of loads, WL the cargo ship can carry before it sinks.


1
Expert's answer
2021-09-25T13:09:18-0400

"F=\rho gV=1020\cdot9.81\cdot6300=63039060\ (N)"


"63039060=mg+WL\to WL=63039060-mg="


"=63039060-500000\cdot9.81=58134060\ (N)" or "m=5926000\ (kg)"